OpenStudy (anonymous):

Write an equation in slope intercept form of the line through point P(-1,-5) with slope -4.

OpenStudy (abbles):

First you need to use point slope form, y - y1 = m(x - x1), and then isolate the y to get it into slope intercept form.

OpenStudy (anonymous):

so if i had to choose between these answers which one would be best?? a. y=-4x-5 b.y+5=-4(x=1) c. y+1=-4(x+5) d.y=-4x-9

OpenStudy (abbles):

Plug the coordinates into the equation and then solve. y - y1 = m(x - x1) y + 5 = -4(x + 1) y + 5 = -4x - 4 y = -4x -9 There you go, that's your answer.

OpenStudy (abbles):

So it would be D.

OpenStudy (anonymous):

thank you so much abbles u saved my life :)

OpenStudy (abbles):

no problem :)

OpenStudy (anonymous):

btw how do you get metals?

OpenStudy (abbles):

Metals can be given when you answer a question, not only by the person that asked but anyone really. So if you help someone they'll probably give you a metal.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

how do i give you a metal?

OpenStudy (abbles):

Someone gave me a metal for this question, I'm not sure if it was you or not but you can check. Do you see the blue button on the right side of each message that says 'best response'? Click that and it will give a metal. :)

OpenStudy (anonymous):

ok i gave u a metal

OpenStudy (abbles):

thanks, I'll give you one too

OpenStudy (anonymous):

thanks

OpenStudy (abbles):

no problem :) hmu if you have any other questions

OpenStudy (anonymous):

what does hmu mean

OpenStudy (abbles):

It means hit me up. You can send me a message or tag me in a question by writing @abbles if you need help with something.

OpenStudy (anonymous):

oh ok where can the perpendicular bisectors of the sides of a right triangle intersect? I. inside the triangle II.on the triangle III. outside the triangle answer choices: a. I,II OR II b. I only c. I or II only d. II only @Abbles

OpenStudy (abbles):

I think the answer is D, II only because the two sides of a triangle would intersect on the triangle itself.

OpenStudy (anonymous):

thanks again for your help

OpenStudy (abbles):

Yep no problem

OpenStudy (anonymous):

i became your fan

OpenStudy (anonymous):

i have another question posted can u help me with it please

OpenStudy (abbles):

I'm looking at it but I can't view the pic for some reason..